Find P(z>-2.575). a. 0.0050 b. 0.6500 c. 0.1950 d. 0.9950

Find P(z>-2.575).

a. 0.0050

b. 0.6500

c. 0.1950

d. 0.9950
Answer #1

Solution :
P(Z > -2.575 ) = 1 - P(Z < -2.575) =
Using standard normal table
P(Z < z) =
To see the z value - 0.2 in the column and 0.575 in the row of the standard normal table the corresponding probability is

= 1 - P(Z < - 2.575 )

= 1 -  0.0050

= 0.9950
Probability = = 0.9950

Option d ) is correct.